OpenStudy (anonymous):

Evaluate (if possible) the sine, cosine, and tangent of the real number. (If not possible, enter IMPOSSIBLE.) t = −π/4

OpenStudy (anonymous):

ok, so convert to degrees \[-\frac{\pi}{4} \text{ }^{c} = -45^{o}\] lets also draw out the trig graphs:  can you see that sin(-x) = -sin(x) and cos(-x) = cos(x) and tan(-x) = -tan(x) ? so if we can work out the trig values of 45 then we can work out values for -45